# Theory of Mind abilities of Large Language Models in Human-Robot Interaction : An Illusion?

### 10 Jan 2024 v1

@@ 4.2.1. Question Design
For answering research question 2, we designed the first question as a binary response question asking the participant a ‘Yes/No’ question on the agent behavior being explicable (and, similarly for the other three parts). The second question was a Multiple Choice Question (MCQ) asking the participants to choose the correct reason behind the expected agent behavior. In the third question, we presented the participants with GPT-4’s response when given the same prompt comprising of the domain description and the first question (see, vignettes in Supplementary for the complete prompts). For the last question in each part, participants had to choose a value on a Likert Scale from 1 to 5, where 1 corresponded to Strongly Disagree, and 5 corresponded to Strongly Agree with the LLM’s reasoning on the expected agent behavior. Our study had a mixed design: Test Domains were varied between-subjects, and response ordering was varied within-subjects for the first two questions.
@@ 4.2.2. Setup & Procedure
The study was setup on the Prolific platform (Palan and Schitter, 2018) where participants were provided with a link directing them to a Google Form to collect their responses. Participants could take the study on any device including a mobile, laptop or a desktop, and did not require any additional setup warranting the use of audio, camera, microphone or any special software. First, they were presented with the data confidentiality statement, followed questions collecting their demographic details, and finally provided with the instructions and overview of the entire study which consisted of the four parts comprising three questions each. The study approximately took 6 minutes to complete, and participants were compensated at the rate of \$9/hour.
@@ 4.2.3. Participants
We recruited 125 participants on Prolific, where participants were required to be between the ages of 18 and 60. We conducted a pilot study on 25 participants spread across each of the five evaluation domains. The final study, refined using the pilot study responses, had a sample size of 100 participants with 20 participants for each of the five domains listed in Section 5, and their demographic details have been shown in Table 1. Most participants fell in the age range of 21 to 34 years, and had an undergraduate degree. Participants reported a mean familiarity of 3.07 and standard deviation of 0.36 to Machine Learning and Artificial Intelligence tools/technologies, on a 7-point Likert Scale with 1 corresponding to being least familiar.
@@ 4.3. Testing LLMs for Robustness
Figure 4. An illustrative view of the evaluation domains used for testing LLM responses for Theory of Mind reasoning on the Perceived Behavior Recognition task. Left to Right: Fetch Robot Domain, Passage Gridworld Domain, Environment Design Domain, Urban Search and Rescue Domain, and Package Delivery Domain.
@@
Inconsistent Belief: An essential component to checking whether mental-model reasoning abilities exist, is by including an additional context which ensures that the correct response can only be computed through semantic reasoning given all the information in the task prompt. Specifically, we include the sentence “The human observer can not see the robot acting in the environment.” in the middle of each prompt. In this case, we modify the options for the first binary response question to also include Can’t Say (the correct answer) along with Yes/No. We also modify the options for the MCQ response question to include the correct reason stating the inability of the human observer to see and understand the agent behavior as given in the rest of the prompt.
Description: Fetch is a robot that can carry objects from one location to another. There are three locations: loc1, loc2, and loc3 where the robot can go. There is a block b1 at location loc1, and the robot is at location loc1 and has its hand empty. Location loc2 is to the left of loc1, and location loc3 is to the right of loc1. Goals: The robot has to take the block b1 to either loc2 OR loc3 (only one of these locations). The human observer can not see the robot acting in the environment. Definition : A partial plan is a part of robot’s behavior, for example a few actions that it takes. Definition : A partial plan is legible if the observer (you) can identify which goal the robot wants to go for. A partial plan A is more legible than another partial plan B if the number of possible goal locations for A is less than B. Plan : In the robot’s partial plan, it picks block b1 and takes one step left. Question 1 : Imagine you are the human observer. Would you find such a partial plan legible? Give your answer as ’Yes’ or ’No’ only.
Description: Fetch is a robot that can carry objects from one location to another. There are three locations: loc1, loc2, and loc3 where the robot can go. There is a block b1 at location loc1, and the robot is at location loc1 and has its hand empty. Location loc2 is to the left of loc1, and location loc3 is to the right of loc1. Goals: The robot has to take the block b1 to either loc2 OR loc3 (only one of these locations). Definition : A partial plan is a part of robot’s behavior, for example a few actions that it takes. Definition : A partial plan is legible if the observer (you) can identify which goal the robot wants to go for. A partial plan A is more legible than another partial plan B if the number of possible goal locations for A is less than B. Plan : In the robot’s partial plan, it picks block b1 and takes one step left. The robot has a label saying ’Not a Legible Agent’. The human observer cannot read. The human observer looks at the label. The human observer has not seen the robot before. Question 1 : Imagine you are the human observer. Would you find such a partial plan legible? Give your answer as ’Yes’, ’No’, or ’Can’t say’ only.
Conviction Test : Ideally, for a system that exhibits mental-model reasoning abilities, we would expect the system to consistently pick the same response (ideally, the correct response) irrespective of the number of times it is queried with the same input. In the case of LLMs (specifically GPT-X family), the temperature (or $\tau$) parameter controls the stochastic nature of the queried responses. It can be varied in the range of 0-2, with 0 corresponding to a least stochastic response and 2 corresponding to the maximum possible stochasticity which could vary the LLM outputs every time it is queried with a given task.
@@ 5. Evaluation Domains
Description: Fetch is a robot that can carry objects from one location to another. There are three locations: loc1, loc2, and loc3 where the robot can go. There is a block b1 at location loc1, and the robot is at location loc1 and has its hand empty. Location loc2 is to the left of loc1, and location loc3 is to the right of loc1. Goals: The robot has to take the block b1 to either loc2 OR loc3 (only one of these locations). Definition : A partial plan is a part of robot’s behavior, for example a few actions that it takes. Definition : A partial plan is legible if the observer (you) can identify which goal the robot wants to go for. A partial plan A is more legible than another partial plan B if the number of possible goal locations for A is less than B. Plan : In the robot’s partial plan, it picks block b1 and takes one step left. Question 1 : Would you find such a partial plan legible? Give your answer as Yes or No only.
@@
Description: Consider a 4x4 square grid with each cell numbered as (row, column), the robot needs to travel from top left cell 1 (1,1) to its goal at bottom right cell 15 (4,3). The human observer (you) expects the robot to take the shortest path by going DOWN 3 steps to row 4 (reach 4, 1), and then RIGHT 2 steps to column 3 (reach 4,3). Constraint: There are blockades in columns 1, 2 and 3, that the human observer (you) do not know of, but the robot does. Definition : Plan Explicability means whether the plan / robot behavior is an expected behavior according to the human observer (you). If you look at the robot behavior and find that some actions are unnecessary or not required, then the behavior is inexplicable. Plan: The robot goes RIGHT 3 steps in row 1 (reach 1,4), goes DOWN 3 steps in column 4 (reach 4,4), and LEFT 1 step in row 4 (4,3). Question 1: Imagine you are the human observer, would you find such a plan explicable? Give your answer as Yes or No only.
Environment Design Gridworld: In contrast to the remaining domains that we consider, the environment design domain tests LLMs on the design choices of a gridworld environment that correspond to an agent’s explicable behavior (and for other three behavior types).
@@
Description: Consider a situation where there is a robot which manages a shipping port, and a human observer (you) who is the supervisor that has sensors or subordinates at the port who provide partial information about the nature of activity being carried out at the port. For instance, when a specific crate is loaded onto the ship, the observer finds out about the identity of the loaded crate. The observer knows the initial inventory at the port, but when new cargo is acquired by the port, the observer’s sensors reveal only that more cargo was received; they do not specify the numbers or identities of the received crates. Initial state: There are packages at the port. Goal: The robot can either pick and acquire a package on the port, or else pick and load it on the ship. Definition: Suppose you think the robot is trying to achieve one out of a set of of potential goals. If the agent’s behavior does not reduce the size of this set, then it is obfuscatory. For example, if you think robot is trying to achieve one of A, B, C. If it shows a behavior (partial plan) but you think it is still trying to achieve any one of A, B, C, then it is obfuscatory. Plan : Suppose the agent picks up the package and holds it between the port and the ship. Question 1 : Would you find such a partial plan obfuscatory? Give your answer as Yes or No only.
USAR: Urban Search And Rescue is a popular domain where the robot and a human observer (commander) perform triage and rescue operations. The robot uses the LLM to query whether a certain plan will be perceived by the commander.
Description : In a typical Urban Search and Rescue (USAR) setting, there is a building with interconnected rooms and hallways. There is a human commander CommX, and a robot agent acting in the environment. Both the agents can move around and pickup/drop-off or handover med-kits to each other. CommX can only interact with med-kits light in weight, but the robot agent can interact with heavy med-kits too. Initial State : There is a medkit located in the middle of the hallway, and there are two paths to a patient room from there (one on the LEFT, and one on the RIGHT). The observer (you) has the top-view of this setting. Goal : Agent has to pickup medkit and take it to the patient room. Definition: A partial plan A is predictable if the observer (you) can identify if there is one possible completion (which may or may not lead to the goal). A partial plan A is more predictable than a partial plan B if the number of possible completions of A is less than B. Plan : In the robot’s partial plan, it picks up the medkit, and takes one step LEFT. Question 1 : Would you find such a partial plan predictable? Give your answer as Yes or No only.
@@ 6. Results
We breakdown our discussion on the results to answer the three research questions, as mentioned in Section 4. In summary, we begin with evaluating lay users performance on our five evaluation domains in 6.1, which validates our evaluation prompts, and also acts as a baseline when comparing LLM performance in 6.2. Finally, we shift our focus to testing the robustness of LLMs with an in-depth analysis of their failure modes in 6.3.
6.1. RQ1 : Lay Users on ToM-PROBE task
@@
Figure 5. Performance accuracy of user study participants across the five domains.
6.2. RQ2: LLMs Performance Alignment with Lay Users
@@ 6.2.1. Performance Comparison :
To answer our second research question (RQ2), we begin with comparing the performance accuracy of GPT-4 and GPT-3.5-turbo with lay users and oracle, as shown in Figure 6. To our expectations, GPT-4 performs better than GPT-3.5-turbo in all but one behavior type, i.e., for predictability where their accuracies are equal.
Figure 6. Performance comparison across the four robot behavior types using Q1 (binary response) results. Each bar represents the number of correct answers to Q1 in the five domains along each behavior type. Note that the lay users’ results have been scaled for a uniform comparison.
@@ 6.2.2. Kolmogorov-Smirnov (KS) Test :
For further evaluating the alignment between LLM responses and lay-users, we perform the KS test between the two populations of (scaled) user responses and GPT-4 responses with the null hypothesis $H_{0}$: the two populations belong to the same distribution. For a KS test with sample size 20, the critical value for significance level 0.01 is 0.356, and for significance level 0.05 is 0.294. We observe that the observed critical value is 0.04, and hence, we accept the null hypothesis. We also run the KS test between GPT-4 responses and the ground truth results, and obtain 0.3 as the observed critical value at which $H_{0}$ can only be accepted for significance level 0.05 and not for significance level 0.01. These tests have been run using the GPT-4 results obtained with the hyperparameter $\tau$ (temperature) set to 0. With $\tau=1$, we obtain the similar results with the observed critical values 0.07 and 0.35 in the two respective cases. Hence, we conclude from these results that GPT-4’s performance aligns with that of lay users as compared to the case of oracles or ground truth.
@@ 6.2.3. Subjective Response Evaluation:
Additionally, for Q3 of in our user study, we record the average scores for human evaluations on LLM descriptive reasoning responses. These ratings are in the [1-5] Likert scale where 1 implies users strongly disgree and at 5 users strongly agree with the GPT-4’s explanation. We find that GPT-4 descriptive responses to vanilla prompt are generally agreed by the users (with average Likert rating of Explicability : 3.47, Legibility : 4.23, Predictability : 4.0, Obfuscation : 4.08) further bolstering the (illusion) belief that LLMs are in fact performing second order ToM (c.f. Figure 13 in the Appendix for more details). For completene we use an automated method of comparing the correct explanation in Q2 and descriptive explanation by GPT-4 by using gpt-3.5 as a semantic text similarity tool (see prompt B.1 in Appendix for more details) in Figure 8, with average ratings of Explicability : 3.4, Legibility : 4, Predictability : 3.8, Obfuscation : 3.4.
@@ 6.3.3. Conviction Test
Figure 9. Performance accuracy of GPT-4 showcasing failure modes in perturbation tests. Each bar represents the number of correct answers to Q1 in the five domains along each behavior type.
@@ 7. Case Study
Extended Research Question 1: How do lay users perform on HRI ToM-PROBE when observing a real robot?
@@ 7.2. Results
We ask the human subjects whether they can read the text on the robot’s head (Fig 1:center or see the robot acting Fig 1:right) as a binary question. 100% users responded they could not read the text and 95% could not see the robot acting which validates our operationalization of the perturbation. We find that the user subject responses and accuracy (90% users correctly identified the behavior-type as legible and 85% users gave correct answer to Q2) are consistent with Fetch-Legibility text-only case which establishes our baseline that human users can answer ToM queries for the constructed vanilla situation. We find that the users are not affected by the Informative Context (85% users correctly identified behavior-type and 95% users chose the correct explanation in Q2) and Inconsistent Belief perturbations (75% correctly opted for “Can’t Say” choice and 25% chose “No” option). The high accuracy values between vanilla case and perturbation case shows that users were not affected by perturbations.

### 17 Jan 2024 v2

@@ 4.2.1. Question Design
For answering RQ2, we designed the first question as a binary Yes/No response on the agent behavior conforming to the given behavior type. The second question was a Multiple Choice Question (MCQ) asking the participants to choose the correct reason behind the agent behavior. In Q3, we presented the participants with GPT-4’s response when given the same prompt comprising of the domain description and the first question (see, vignettes in Supplementary). For the last question, participants answered on a Likert Scale from 1 (Strongly Disagree) to 5 (Strongly Agree) with the LLM’s reasoning on the agent behavior. Our study had a mixed design: Test Domains were varied between-subjects, and response ordering was varied within-subjects for the first two questions.
@@ 4.2.2. Setup & Procedure
The study used the Prolific platform (Palan and Schitter, 2018), directing participants to a Google Form for responses. Participants could use any device without requiring speaker, camera, microphone, or special software for the study. First, they were presented with the data confidentiality statement, followed questions collecting their demographic details, and finally provided with the instructions and overview of the entire study which consisted of the four parts comprising three questions each. The study approximately took 6 minutes to complete, and participants were compensated at the rate of \$9/hour.
@@ 4.2.3. Participants
We recruited 125 participants (between the ages of 18 and 60). We conducted a pilot study on 25 participants spread across each of the five evaluation domains. The final study, refined using the pilot study responses, had a sample size of 100 participants with 20 participants for each of the five domains listed in Section 5 (see Table 1). Most participants fell in the age range of 21 to 34 years, and had an undergraduate degree. Participants reported a mean familiarity of 3.07 and standard deviation of 0.36 to Machine Learning and Artificial Intelligence tools/technologies, on a 7-point Likert Scale with 1 corresponding to being least familiar.
@@ 4.3. Testing LLMs for Robustness
Figure 3. An illustrative view of the domains used for testing LLM for Theory of Mind reasoning on the Perceived Behavior Recognition task. Left to Right: Fetch Robot Domain, Passage Gridworld, Environment Design, Urban Search and Rescue, and Package Delivery.
@@
Description: Fetch is a robot that can carry objects from one location to another. There are three locations: loc1, loc2, and loc3 where the robot can go. There is a block b1 at location loc1, and the robot is at location loc1 and has its hand empty. Location loc2 is to the left of loc1, and location loc3 is to the right of loc1. Goals: The robot has to take the block b1 to either loc2 OR loc3 (only one of these locations). Definition : A partial plan is a part of robot’s behavior, for example a few actions that it takes. Definition : A partial plan is legible if the observer (you) can identify which goal the robot wants to go for. A partial plan A is more legible than another partial plan B if the number of possible goal locations for A is less than B. Plan : In the robot’s partial plan, it picks block b1 and takes one step left. The robot has a label saying ’Not a Legible Agent’. The human observer cannot read. The human observer looks at the label. The human observer has not seen the robot before. Question 1 : Imagine you are the human observer. Would you find such a partial plan legible? Give your answer as ’Yes’ or ’No’ only.
Description: Fetch is a robot that can carry objects from one location to another. There are three locations: loc1, loc2, and loc3 where the robot can go. There is a block b1 at location loc1, and the robot is at location loc1 and has its hand empty. Location loc2 is to the left of loc1, and location loc3 is to the right of loc1. Goals: The robot has to take the block b1 to either loc2 OR loc3 (only one of these locations). The human observer can not see the robot acting in the environment. Definition : A partial plan is a part of robot’s behavior, for example a few actions that it takes. Definition : A partial plan is legible if the observer (you) can identify which goal the robot wants to go for. A partial plan A is more legible than another partial plan B if the number of possible goal locations for A is less than B. Plan : In the robot’s partial plan, it picks block b1 and takes one step left. Question 1 : Imagine you are the human observer. Would you find such a partial plan legible? Give your answer as ’Yes’, ’No’, or ’Can’t say’ only.
@@
Inconsistent Belief: An essential component to checking whether mental-model reasoning abilities exist, is by including an additional context which ensures that the correct response can only be computed through semantic reasoning. Specifically, we include the sentence “The human observer can not see the robot acting in the environment.” in each prompt. We modify the options for the first binary response question to include Can’t Say (the correct answer) along with Yes/No. We also modify Q2 to include the correct reason stating the inability of the human observer to see and understand the agent behavior.
Conviction Test : For a system that exhibits mental-model reasoning abilities, we would expect the system to consistently pick the same response (ideally, the correct response) regardless of the repeated queries. In the case of GPT-X family, the temperature (or $\tau$) parameter controls the stochastic nature of the queried responses. It can be varied in the range of 0 (least)-2 (most) stochasticity influencing the LLM outputs on each iteration.
@@ 5. Evaluation Domains
Description: Fetch is a robot that can carry objects from one location to another. There are three locations: loc1, loc2, and loc3 where the robot can go. There is a block b1 at location loc1, and the robot is at location loc1 and has its hand empty. Location loc2 is to the left of loc1, and location loc3 is to the right of loc1. Goals: The robot has to take the block b1 to either loc2 OR loc3 (only one of these locations). Definition : A partial plan is a part of robot’s behavior, for example a few actions that it takes. Definition : A partial plan is legible if the observer (you) can identify which goal the robot wants to go for. A partial plan A is more legible than another partial plan B if the number of possible goal locations for A is less than B. Plan : In the robot’s partial plan, it picks block b1 and takes one step left. Question 1 : Would you find such a partial plan legible? Give your answer as ‘Yes’ or ‘No’ only.
@@
Description: Consider a 4x4 square grid with each cell numbered as (row, column), the robot needs to travel from top left cell 1 (1,1) to its goal at bottom right cell 15 (4,3). The human observer (you) expects the robot to take the shortest path by going DOWN 3 steps to row 4 (reach 4, 1), and then RIGHT 2 steps to column 3 (reach 4,3). Constraint: There are blockades in columns 1, 2 and 3, that the human observer (you) do not know of, but the robot does. Definition : Plan Explicability means whether the plan / robot behavior is an expected behavior according to the human observer (you). If you look at the robot behavior and find that some actions are unnecessary or not required, then the behavior is inexplicable. Plan: The robot goes RIGHT 3 steps in row 1 (reach 1,4), goes DOWN 3 steps in column 4 (reach 4,4), and LEFT 1 step in row 4 (4,3). Question 1: Imagine you are the human observer, would you find such a plan explicable? Give your answer as ‘Yes’ or ‘No’ only.
Environment Design Gridworld: The environment design domain tests LLMs on the design choices of a gridworld environment that correspond to an agent’s explicable behavior (and for other three behavior types).
@@
Description: Consider a situation where there is a robot which manages a shipping port, and a human observer (you) who is the supervisor that has sensors or subordinates at the port who provide partial information about the nature of activity being carried out at the port. For instance, when a specific crate is loaded onto the ship, the observer finds out about the identity of the loaded crate. The observer knows the initial inventory at the port, but when new cargo is acquired by the port, the observer’s sensors reveal only that more cargo was received; they do not specify the numbers or identities of the received crates. Initial state: There are packages at the port. Goal: The robot can either pick and acquire a package on the port, or else pick and load it on the ship. Definition: Suppose you think the robot is trying to achieve one out of a set of of potential goals. If the agent’s behavior does not reduce the size of this set, then it is obfuscatory. For example, if you think robot is trying to achieve one of A, B, C. If it shows a behavior (partial plan) but you think it is still trying to achieve any one of A, B, C, then it is obfuscatory. Plan : Suppose the agent picks up the package and holds it between the port and the ship. Question 1 : Would you find such a partial plan obfuscatory? Give your answer as ‘Yes’ or ‘No’ only.
USAR: Urban Search And Rescue is a popular domain where the robot and a human observer (commander) perform triage and rescue operations. The robot uses the LLM to query how a certain plan will be perceived by the commander.
Description : In a typical Urban Search and Rescue (USAR) setting, there is a building with interconnected rooms and hallways. There is a human commander CommX, and a robot agent acting in the environment. Both the agents can move around and pickup/drop-off or handover med-kits to each other. CommX can only interact with med-kits light in weight, but the robot agent can interact with heavy med-kits too. Initial State : There is a medkit located in the middle of the hallway, and there are two paths to a patient room from there (one on the LEFT, and one on the RIGHT). The observer (you) has the top-view of this setting. Goal : Agent has to pickup medkit and take it to the patient room. Definition: A partial plan A is predictable if the observer (you) can identify if there is one possible completion (which may or may not lead to the goal). A partial plan A is more predictable than a partial plan B if the number of possible completions of A is less than B. Plan : In the robot’s partial plan, it picks up the medkit, and takes one step LEFT. Question 1 : Would you find such a partial plan predictable? Give your answer as ‘Yes’ or ‘No’ only.
@@ 6. Results
We breakdown our discussion on the results to answer the three research questions, as mentioned in Section 4. In summary, we begin with evaluating human subjects performance on our five evaluation domains in 6.1, which validates our evaluation prompts, and also acts as a baseline when comparing LLM performance in 6.2. Finally, we shift our focus to testing the robustness of LLMs with an in-depth analysis of their failure modes in 6.3.
6.1. RQ1 : Human Subjects on ToM-PROBE task
@@
Figure 4. Human subject performance across five domains.
6.2. RQ2: LLMs Performance Alignment
@@ 6.2.1. Performance Comparison :
To answer our second research question (RQ2), we begin with comparing the performance accuracy of GPT-4 and GPT-3.5-turbo with human subjects and oracle, as shown in Figure 5. To our expectations, GPT-4 performs better than GPT-3.5-turbo in all but one behavior type, i.e., for predictability where their accuracies are equal.
Figure 5. Performance on Q1 across five domains along four robot behavior types on Q1 (binary response). Human subjects’ results have been scaled for a uniform comparison.
Figure 7. Performance accuracy of GPT-4 showcasing failure modes in perturbation tests. Each bar denotes the correct answers for Q1 across five domains along behavior types.
@@ 6.2.2. Kolmogorov-Smirnov (KS) Test :
For further evaluating the alignment between LLM responses and human subjects, we perform the KS test between the two populations of (scaled) user responses and GPT-4 responses with the null hypothesis $H_{0}$: the two populations belong to the same distribution. For a KS test with sample size 20, the critical value for significance level 0.01 is 0.356, and for significance level 0.05 is 0.294. We observe that the observed critical value is 0.04, and hence, we accept the null hypothesis. We also run the KS test between GPT-4 responses and the ground truth results, and obtain 0.3 as the observed critical value at which $H_{0}$ can only be accepted for significance level 0.05 and not for significance level 0.01. These tests have been run using the GPT-4 results obtained with the hyperparameter $\tau$ (temperature) set to 0. With $\tau=1$, we obtain the similar results with the observed critical values 0.07 and 0.35 in the two respective cases. Hence, we conclude from these results that GPT-4’s performance aligns with that of human subjects as compared to the case of oracles or ground truth.
@@ 6.2.3. Subjective Response Evaluation:
Additionally, for Q3 of in our user study, we record the average scores for human evaluations on LLM descriptive reasoning responses. These ratings are in the [1-5] Likert scale where 1 implies users strongly disgree and at 5 users strongly agree with the GPT-4’s explanation. We find that GPT-4 descriptive responses to vanilla prompt are generally agreed by the users (with average Likert rating of Explicability : 3.47, Legibility : 4.23, Predictability : 4.0, Obfuscation : 4.08) further bolstering the (illusion) belief that LLMs are in fact performing second order ToM (c.f. Figure 12 in the Appendix for more details). For completeness, we use an automated method of comparing the correct explanation in Q2 and descriptive explanation by GPT-4 by using gpt-3.5 as a semantic text similarity tool (see prompt A.1 in Appendix for more details) in Figure 6, with average ratings of Explicability : 3.4, Legibility : 4, Predictability : 3.8, Obfuscation : 3.4.
@@ 7. Case Study
Extended Research Question 1: How do human subjects perform on HRI ToM-PROBE when observing a real robot?
@@ 7.2. Results
We ask the human subjects whether they can read the text on the robot’s head (Fig 1:center) or see the robot acting (Fig 1:right) as a binary question. 100% users responded they could not read the text and 95% could not see the robot acting which validates our operationalization of the perturbation. GPT4-Vision (OpenAI, 2023a) is also fed with image frames obtained via stratified sampling (step size $k=\{6,12,24,30\}$) from the video along with the text prompt. Across all sampling step sizes, the LLM always responds that it cannot read the text and cannot see the robot acting in the respective cases. We find that the user subject responses and accuracy (90% users correctly identified the behavior-type as legible and 85% users gave correct answer to Q2) are consistent with Fetch-Legibility text-only case which establishes our baseline that human users can answer ToM queries for the constructed vanilla situation. We find that the users are not affected by the Uninformative Context (85% users correctly identified behavior-type and 95% users chose the correct explanation in Q2) and Inconsistent Belief perturbations (75% correctly opted for “Can’t Say” choice and 25% chose “No” option). The high accuracy values between vanilla case and perturbation case shows that users were not affected by perturbations. GPT4-V correctly identifies the behavior-type along with the correct reasoning in the vanilla case, however, it fails in the Inconsistent Belief test and the Conviction test. We note that, while GPT4-V answers that it cannot see the robot acting, it still fails Q1 which also reflects inconsistent reasoning abilities. More details can be found in Section A.3.